This is from a worksheet: In a chemistry class, 6 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% solution is needed?
Answer by stanbon(75887)

In a chemistry class, 6 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% solution is needed?
----------------
Amt of 4% solution = 6 L ; Amt of active ingredient = 0.04*6 = 0.24 L
-------------
Amt of 10% solution = x L ; Amt of active ingredient = 0.10x L
-----------------
Amto of 6% mixture = "6+x" L; Amt of active = 0.06(6+x)=0.36+0.06x L
--------------
EQUATION:
active + active = active
0.24 + 0.10x = 0.36+0.06x
0.04x = 0.12
x = 3 L (Amt of 10% solution that must be added)
============
Cheers,
Stan H.